Joe is six years older than his brother. In one more year he will be twice as old. Find their ages now.
I used the formula x+6 = 2(x+1) and got x = 4

Answer by stanbon(75887)

Joe is six years older than his brother. In one more year he will be twice as old. Find their ages now.
I used the formula x+6 = 2(x+1) and got x = 4
---------------
Equations:
j = b + 6
j+1 = 2(b+1)
-----
Substitute for "j" and solve for "b":
b+6+1 = 2b + 2
b = 5 (his brother's age now)
-----
Solve for "j":
j = b + 6
j = 11 yrs (joe's age now)
---------------------------
Cheers,
Stan H.
====================